Mary Alice just won the lottery and is trying to decide between the options of receiving the annual cash flow payment option of $300,000 per year for 25 years beginning today, or receiving one lump-sum amount today. Mary Alice can earn 5% investing this money. At what lump-sum payment amount would she be indifferent between the two alternatives? (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Multiple Choice

  • $4,739,592

  • $4,439,592

  • $7,500,000

  • $4,228,182
Answer #1

Answer: $ 4,439,592

The present value of the First option

300,000 per Every year, for 25 years

= 300,000+(PVA at 5% for 25 years ) *300,000= 300,000+(300000*13.79864)

= $4,439,592 (Answer)

indifference between Payng $ 4,439,592 todays or 300,000 for 25 years are sames